# SolarGPT-QA: A Domain-Adaptive Large Language Model for Educational Question Answering in Space Weather and Heliophysics

### v3

@@ Abstract
Solar activity, including solar flares, coronal mass ejections (CMEs), and geomagnetic storms, can significantly impact satellites, aviation, power grids, data centers, and space missions. Extreme solar events can cause substantial economic damage with limited advance warning, underscoring the importance of early-warning systems, accurate forecasting, and effective education in space science. Although large language models (LLMs) perform well on general tasks, they often lack domain-specific knowledge and pedagogical capability to clearly explain complex space science concepts. We introduce SolarGPT-QA, a question answering system based on a domain-adapted large language model built on the LLaMA-3 base model. The model is trained using scientific literature and large-scale question-answer data generated with GPT-4 and refined using Grok-3 in a student-friendly storytelling style. Human pairwise evaluations show that SolarGPT-QA outperforms general-purpose models in zero-shot settings and achieves competitive performance compared to instruction-tuned models for educational explanations in space weather and heliophysics. A small pilot student comprehension study further suggests improved clarity and accessibility of the generated explanations. Ablation experiments indicate that combining domain-adaptive pretraining with pedagogical fine-tuning is important for balancing scientific accuracy and educational effectiveness. This work represents an initial step toward a broader SolarGPT framework for space science education and forecasting.
I Introduction
@@
Heliophysics provides a system-level understanding of the Sun and its influence throughout the heliosphere [39]. Within this broader heliophysical context, space weather science investigates the interconnected and multiscale interactions among the Sun, the solar wind, interplanetary space, and planetary environments, and examines how solar activity influences Earth’s space environment and technological systems [50]. Solar phenomena including solar flares (sudden bursts of electromagnetic radiation from the Sun) [50, 49], coronal mass ejections (CMEs) (large eruptions of magnetized plasma from the solar corona) [64], and solar energetic particles (SEPs) (high-energy charged particles accelerated by flares and CMEs) [47, 3] can trigger geomagnetic storms [23], that propagate through the near-Earth environment, disrupting satellite operations, communication networks, power grids, and data centers, while exposing astronauts and high-altitude aviation to elevated radiation hazards. For instance, the infamous ”Carrington Event” in 1859 [37] was the most intense geomagnetic storm on record, triggered by an extreme solar flare and CME, while the Quebec blackout in 1989 [18] vividly demonstrated the susceptibility of modern technological systems to space weather disturbances. Such disruptions have resulted in economic losses of billions of dollars, and an event of similar scale today could cause even greater societal and technological consequences in our technology-driven society [6, 3]. Additionally, as modern society becomes increasingly reliant on space-based technologies, including large satellite constellations such as Starlink [43], and as private space enterprises like SpaceX and Blue Origin pursue future space-based data centers and exploration missions [58], understanding space weather has emerged as a critical scientific, technological, and educational priority, with relevance extending as early as the high-school level, bringing an additional dimension to its importance.
@@
Large language models have become powerful tools for understanding and explaining complex information, with growing impact across many scientific fields [4]. They can answer questions, summarize research, and support learning. However, their use in space weather and heliophysics is still limited, largely because these areas rely on highly specialized literature and concepts that are rarely covered in general training data. One key challenge is that most space weather knowledge exclusively resides in technical papers and mission reports, which means general models often miss key terms, cause–effect links, or physical limits. Another issue is that teaching space science, particularly to K–12 students or early undergraduates, requires clear language, analogies, and a storytelling approach. General models are not built to balance accuracy that is tailored to the appropriate age.
@@
To fill this gap, we introduce SolarGPT-QA, a domain-adapted model designed for educational question answering in space weather and heliophysics for K–12 learners. The model is trained on a carefully curated corpus of high-impact, peer-reviewed research articles published since 2010, aligned with the Solar Dynamics Observatory (SDO) era of modern flare and CME research. These sources are further refined for relevance to solar eruptions and used to generate a large collection of synthetic question–answer pairs. Rather than targeting expert-level analysis, SolarGPT-QA focuses on helping students grasp core concepts through clear, accessible explanations to questions such as how solar flares form or why the Sun can disrupt technology on Earth.
@@
II Related Work
II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [22]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [32, 42, 24].Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [63, 67], but their operational use is limited by computational cost and incomplete boundary conditions [66]. Furthermore, although significant progress has been made in understanding solar magnetic field dynamics, a quantitative theory that rigorously connects the accumulation and restructuring of magnetic flux to the onset of extreme solar events remains lacking in physics-driven approaches, leaving the key mechanisms governing flare and coronal mass ejection initiation still not fully understood [1]. This gap highlights the importance of data-driven approaches and machine learning methods for improving the prediction and characterization of extreme solar events.
@@ II-A Forecasting of Solar Flares, CMEs, and SEPs
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [25] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters. Progress in this area has been supported by the development of benchmark datasets. The Space Weather Analytics for Solar Flares (SWAN-SF) multivariate time series dataset integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [2]. Machine learning-based models in this domain include classical approaches such as logistic regression [52], support vector machines [7], and decision trees [35], as well as deep learning architectures for temporal sequence modeling, including LSTMs [38], contrastive learning approaches [59, 60], graph-based methods [62], and transformer frameworks [61]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [20, 26].
Many efforts have applied deep learning on solar imagery, including magnetograms and extreme ultraviolet observations, to capture spatial precursors to flares and CMEs, while multimodal fusion of time series and images further enhances forecasting accuracy [41, 68, 27]. Despite these advances, most operational systems predict flares, CMEs, and SEPs independently and provide limited explanatory context, functioning largely as black boxes [17, 44, 45]. These limitations motivate unified, multimodal, and explainable forecasting frameworks that build on prior advances in time series modeling, dataset development, and interpretability.
II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve state-of-the-art performance across wide range of the tasks and model scales [8, 16, 56, 57, 13, 12, 11, 10, 9]. There has been subsequent work and exploration for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [46, 70, 55, 48, 69]. DAPT and fine-tuning are common techniques to improve the LLM for specific datasets or domains. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [30, 28, 21]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [65, 40, 53, 15].
Beyond general-purpose models, large language models have increasingly been adapted to specialized scientific domains, including medicine, biology, Earth science, and environmental sciences. In the medical domain, Med-PaLM [51] demonstrates that large-scale language models can encode clinical knowledge and perform competitively on medical question answering through instruction tuning and expert evaluation. Generalist medical foundation models, such as those proposed by Moor et al. [36], extend this approach by enabling a single model to support a wide range of clinical reasoning tasks. MedGPT [29] leverages electronic health records and named entity recognition to predict future medical events, while BioGPT [34] is pre-trained on large-scale biomedical literature to improve biomedical text generation and knowledge extraction.
Geneformer [54] is a life science domain-adapted language model that applies transformer-based language modeling to single-cell transcriptomic data, enabling transfer learning for gene network inference and cellular state prediction. Similarly, large-scale protein language models demonstrate that language modeling techniques can predict atomic-level protein structures directly from amino acid sequences [33].
More recently, scientific LLMs have been introduced for Earth, ocean, and geoscience applications. GeoGPT [19] represents one of the first large language models explicitly designed for geoscience tasks. OceanGPT [5] is trained on marine and oceanographic literature, demonstrating superior performance and preliminary embodied intelligence for diverse ocean science and engineering tasks compared to general-purpose LLMs. CLLMate [31] focuses on climate-related applications, integrating climate science literature and data-driven knowledge to support climate analysis, interpretation, and decision support. Tele-Knowledge Pretraining [14] incorporates structured technical knowledge into language models for fault diagnosis and engineering analysis.
Despite these advances, most existing scientific LLMs primarily target expert-level analysis and professional decision-making, often producing technically dense and jargon-heavy outputs. In contrast, our work focuses on domain-adapted language modeling for educational scientific question answering, emphasizing conceptual clarity, accessibility, and age-appropriate explanations. Rather than supporting expert users, SolarGPT-QA is explicitly designed to facilitate learning and understanding in complex scientific domains such as space weather and heliophysics.
III SolarGPT-QA
@@
SolarGPT-QA is a domain-adapted large language model designed specifically for educational question answering in space weather and heliophysics, with an emphasis on K–12–friendly explanations. The methodology consists of four main stages: (1) domain corpus construction and preprocessing, (2) domain-adaptive pretraining (DAPT) using LoRA, (3) educational question–answer fine-tuning, and (4) inference-time constraints to promote clarity and reduce hallucinations. An overview of the system architecture is shown in Fig. 2. Let $\mathcal{D}_{H}$ denote a corpus of heliophysics literature, and $\mathcal{D}_{QA}$ denote a supervised question–answer dataset designed for K–12 learners. Our objective is to learn a conditional language model
@@
III-A Domain Corpus Construction and Normalization
@@
III-B Domain-Adaptive Pretraining with Low-Rank Adaptation
We initialize SolarGPT-QA from a pretrained causal language model, Meta-LLaMA-3-8B, and adapt it to the heliophysics domain using domain-adaptive pretraining (DAPT) combined with Low-Rank Adaptation (LoRA) for parameter efficiency. As defined in Section III-A, the heliophysics corpus is processed as
$DH\rightarrow\mathcal{T}$
@@ III-B Domain-Adaptive Pretraining with Low-Rank Adaptation
To enable parameter-efficient adaptation, we employ Low-Rank Adaptation (LoRA). For a transformer weight matrix
$W\in\mathbb{R}^{d\times k},$
LoRA applies a low-rank update
@@
where
$\Delta W=BA.$ (6)
The overall training workflow is summarized in Algorithm 1. Only the low-rank matrices $A$ and $B$ are trainable, while the base model parameters remain frozen. This significantly reduces the number of optimized parameters, and enables efficient domain adaptation without sacrificing the general linguistic capabilities of the pretrained model. Specific LoRA configuration values are provided in Section IV.
III-C Fine-Tuning on QA dataset
@@
where $\mathcal{L}_{\text{DAPT}}$ denotes the domain-adaptive pretraining loss and $\mathcal{L}_{QA}$ denotes the educational question–answer fine-tuning loss, with $\lambda$ controlling the emphasis on pedagogical alignment. In practice, these objectives are optimized sequentially rather than jointly: the model is first adapted to heliophysics literature using $\mathcal{L}_{\text{DAPT}}$, and the resulting DAPT-adapted checkpoint is subsequently fine-tuned using $\mathcal{L}_{QA}$ to produce the final SolarGPT-QA model. The combined objective summarizes the cumulative learning goal across training stages, rather than a single joint optimization.
Unlike prior scientific language models optimized for expert-level analysis, SolarGPT-QA is explicitly designed for educational deployment. By combining domain-adaptive pretraining, parameter-efficient optimization, and pedagogically aligned fine-tuning with constrained decoding, the model bridges expert heliophysics knowledge and accessible K–12 instruction.
III-D Inference-Time Hallucination Mitigation
In addition to domain-adaptive pretraining and educational fine-tuning, SolarGPT-QA applies inference-time decoding constraints to mitigate hallucinated. These constraints regulate generation behavior rather than model parameters, promoting concise, well-formed, and locally grounded responses suitable for educational settings. Controlled stochastic decoding is used to suppress low-probability and repetitive continuations, while generation length and topic drift are limited through a custom sentence-boundary stopping criterion. Post-generation filtering further removes incomplete or syntactically ill-formed sentence fragments, reducing speculative extrapolation beyond the scope of the input question.
IV Experiments
IV-A Training Details and Compute
Domain-adaptive pretraining (DAPT) and QA fine-tuning were performed using LoRA on Meta-LLaMA-3-8B with 8-bit weight loading and FP16 training. LoRA adapters were applied to the attention q_proj and v_proj modules with rank r=16r{=}16, scaling $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training used the AdamW optimizer with a learning rate of $2\times 10^{-4}$, linear warmup of 100 steps, and a maximum sequence length of 512 tokens. The per-device batch size was 1 with gradient accumulation of 8, resulting in an effective batch size of 8 sequences per optimization step. Both DAPT and QA fine-tuning were run for 3 epochs.

### v4

@@ Abstract
Solar activity, including solar flares, coronal mass ejections (CMEs), and geomagnetic storms can significantly impact satellites, aviation, power grids, data centers, and space missions. Extreme solar events can cause substantial economic damage with limited advance warning, underscoring the importance of early warning systems, accurate forecasting, and effective education in space science. Although large language models (LLMs) perform well on general tasks, they often lack domain specific knowledge and pedagogical capability to clearly explain complex space science concepts. We introduce SolarGPT-QA, a question answering system based on a domain adapted large language model built on the LLaMA-3 base model. The model is trained using scientific literature and large scale question and answer data generated with GPT-4 and refined using Grok-3 in a student friendly storytelling style. To evaluate response quality, we employ an LLM-as-judge evaluation framework, where a strong reference model assesses generated answers using structured criteria including scientific accuracy, clarity, completeness, and pedagogical effectiveness. Results show that SolarGPT-QA performs strongly relative to general purpose models in zero shot settings and achieves competitive performance compared to instruction tuned models for educational explanations in space weather and heliophysics. Ablation studies indicate that combining domain adaptive pretraining with fine tuning is important for balancing scientific accuracy and educational effectiveness.
1 Introduction
@@
Heliophysics studies the Sun and its interactions with the surrounding heliosphere, including the processes that drive space weather and influence Earth’s technological environment [nasa_heliophysics]. Within this broader heliophysical context, space weather science investigates the interconnected and multiscale interactions among the Sun, the solar wind, interplanetary space, and planetary environments, and examines how solar activity influences Earth’s space environment and technological systems [schwenn2006space]. Solar activity can trigger geomagnetic storms [gonzalez1994geomagnetic]. Key drivers include solar flares, coronal mass ejections (CMEs), and solar energetic particles (SEPs). These disturbances propagate through the near-Earth environment and can disrupt satellite operations, communication networks, power grids, and data centers. They can also expose astronauts and high-altitude aviation to elevated radiation hazards. For instance, the infamous Carrington Event in 1859 [muller2014carrington] was the most intense geomagnetic storm on record, triggered by an extreme solar flare and CME, while the Quebec blackout in 1989 vividly demonstrated the susceptibility of modern technological systems to space weather disturbances. Such disruptions have resulted in economic losses of billions of dollars, and an event of similar scale today could cause even greater societal and technological consequences in our technology-driven society [blong2021four].
@@
Large language models have become powerful tools for understanding and explaining complex information, with growing impact across many scientific fields [bharathi2024analysis]. They can answer questions, summarize research, and support learning through textual explanations. However, their use in space weather and heliophysics is still limited, largely because these areas rely on highly specialized literature and concepts that are rarely covered in general training data. One key challenge is that most space weather knowledge exclusively resides in journal articles, technical papers, and mission reports, which means general models often miss key terms, cause–effect links, or physical limits. Another issue is that teaching space science, particularly to K–12 students or early undergraduates, requires clear language, analogies, and a storytelling approach. General models are not built to balance accuracy that is tailored to the appropriate age.
@@
To fill this gap, we introduce SolarGPT-QA, a domain-adapted model designed for educational question answering in space weather and heliophysics for K–12 learners. Fig. 1 presents an overview of the SolarGPT-QA educational question–answering framework. The model is trained on a carefully curated corpus of high-impact, peer-reviewed research articles published since 2010, aligned with the Solar Dynamics Observatory (SDO) era of modern flare and CME research. These sources are further refined for relevance to solar eruptions and used to generate a large collection of synthetic question–answer pairs. Rather than targeting expert-level analysis, SolarGPT-QA focuses on helping students grasp core concepts through clear, accessible explanations to questions such as how solar flares form or why the Sun can disrupt technology on Earth.
@@
2 Related Work
2.1 Forecasting of Solar Flares, CMEs, and SEPs
@@
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [georgoulis2024prediction]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [lin2003theories]. Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [wang2013mhd], but their operational use is limited by computational cost and incomplete boundary conditions [whitman2023sepmodels]. Furthermore, although significant progress has been made in understanding solar magnetic field dynamics, a quantitative theory that rigorously connects the accumulation and restructuring of magnetic flux to the onset of extreme solar events remains lacking in physics-driven approaches, leaving the key mechanisms governing flare and coronal mass ejection initiation still not fully understood [ahmadzadeh2021train]. This gap highlights the importance of data-driven approaches and machine learning methods for improving the prediction and characterization of extreme solar events. The launch of NASA’s Solar Dynamics Observatory (SDO) in 2010 marked a major advance in this area by providing continuous, high-resolution observations of the Sun. Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Hamdi et al. [hamdi2017timeseries] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters. Progress in this area has been supported by the development of benchmark datasets. The Space Weather Analytics for Solar Flares (SWAN-SF) multivariate time series dataset integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [angryk2020swansf]. Machine learning-based models in this domain include classical approaches such as logistic regression [song2009statistical], support vector machines [bobra2015solar], and decision trees [ma2017decisiontrees], as well as deep learning architectures for temporal sequence modeling, including LSTMs [muzaheed2021sequence], contrastive learning approaches, graph-based methods, and transformer frameworks [vural2024contrastive, vural2025global]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [eskandarinasab2024preprocessing]. Many efforts have applied deep learning on solar imagery, including magnetograms and extreme ultraviolet observations, to capture spatial precursors to flares and CMEs, while multimodal fusion of time series and images further enhances forecasting accuracy [park2018cnn, hosseinzadeh2025multimodal]. Despite these advances, most operational systems predict flares, CMEs, and SEPs independently and provide limited explanatory context, functioning largely as black boxes [collado2019ccmc, petrenko2023ccmc, petrenko2024barriers]. These limitations motivate unified, multimodal, and explainable forecasting frameworks that build on prior advances in time series modeling, dataset development, and interpretability.
2.2 LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years and can achieve state-of-the-art performance across a wide range of tasks and model scales [brown2020language, Chapagain2025AdvancingMinority]. Subsequent work has explored the scaling behavior of LLMs, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [rae2021scaling, zhang2022opt]. DAPT and fine-tuning are common techniques for improving LLMs for specific datasets or domains. Similarly, retrieval-augmented generation (RAG) has been widely used to improve factual grounding by integrating external knowledge sources into language models [lewis2020rag]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [wei2022finetuned]. Beyond general-purpose models, large language models have increasingly been adapted to specialized scientific domains, including medicine, biology, Earth science, and environmental sciences. In the medical domain, Med-PaLM [singhal2022large] demonstrates that large-scale language models can encode clinical knowledge and perform competitively on medical question answering through instruction tuning and expert evaluation. Generalist medical foundation models, such as those proposed by Moor et al. [moor2023med], extend this approach by enabling a single model to support a wide range of clinical reasoning tasks. MedGPT [kraljevic2021medgpt] leverages electronic health records and named entity recognition to predict future medical events, while BioGPT [luo2022biogpt] is pre-trained on large-scale biomedical literature to improve biomedical text generation and knowledge extraction. Geneformer [theodoris2023geneformer] is a life science domain-adapted language model that applies transformer-based language modeling to single-cell transcriptomic data, enabling transfer learning for gene network inference and cellular state prediction. Similarly, large-scale protein language models demonstrate that language modeling techniques can predict atomic-level protein structures directly from amino acid sequences [lin2023protein]. More recently, scientific LLMs have been introduced for Earth, ocean, and geoscience applications. GeoGPT [deng2023geoscience] represents one of the first large language models explicitly designed for geoscience tasks. OceanGPT [bi2024oceangpt] is trained on marine and oceanographic literature, demonstrating superior performance and preliminary embodied intelligence for diverse ocean science and engineering tasks compared to general-purpose LLMs. CLLMate [li2025cllmate] focuses on climate-related applications, integrating climate science literature and data-driven knowledge to support climate analysis, interpretation, and decision support. Tele-Knowledge Pretraining [chen2023tele] incorporates structured technical knowledge into language models for fault diagnosis and engineering analysis. Despite these advances, most existing scientific LLMs primarily target expert-level analysis and professional decision-making, often producing technically dense and jargon-heavy outputs. In contrast, our work focuses on domain-adapted language modeling for educational scientific question answering, emphasizing conceptual clarity, accessibility, and age-appropriate explanations. Rather than supporting expert users, SolarGPT-QA is explicitly designed to facilitate learning and understanding in complex scientific domains such as space weather and heliophysics.
3 SolarGPT-QA
@@
SolarGPT-QA is a domain-adapted large language model designed specifically for educational question answering in space weather and heliophysics, with an emphasis on K–12–friendly explanations. The methodology consists of four main stages: (1) domain corpus construction and preprocessing, (2) domain-adaptive pretraining (DAPT) using LoRA, (3) educational question–answer fine-tuning, and (4) inference-time constraints to promote clarity and reduce hallucinations. An overview of the system architecture is shown in Fig. 2. Let $\mathcal{D}_{H}$ denote a corpus of heliophysics literature, and $\mathcal{D}_{QA}$ denote a supervised QA dataset designed for K–12 learners. Our objective is to learn a conditional language model
@@
3.1 Domain Corpus Construction and Normalization
@@
3.2 Domain-Adaptive Pretraining with Low-Rank Adaptation
We initialize SolarGPT-QA from a pretrained causal language model, Meta-LLaMA-3-8B, and adapt it to the heliophysics domain using domain-adaptive pretraining (DAPT) [gururangan2020don] combined with Low-Rank Adaptation (LoRA) [hu2022lora] for parameter efficiency. As defined in Section 3.1, the heliophysics corpus is processed as
$\mathcal{D}_{H}\rightarrow\mathcal{T}$
@@ 3.2 Domain-Adaptive Pretraining with Low-Rank Adaptation
For a transformer weight matrix, $W\in\mathbb{R}^{d\times k}$, LoRA applies a low-rank update
@@
Here, $\Delta W$ is the learned adaptation applied to the frozen pretrained weight matrix $W$. LoRA parameterizes this update as the product of two low-rank matrices, $\Delta W=BA$, where $A\in\mathbb{R}^{r\times k}$ projects the original representation into a lower-dimensional space and $B\in\mathbb{R}^{d\times r}$ projects it back to the original dimension. The rank $r\ll\min(d,k)$ controls the number of trainable parameters and enables parameter-efficient adaptation. The overall training workflow is summarized in Algorithm 1, and the LoRA configuration values are provided in Section 4.
3.3 Fine-Tuning on QA dataset
@@
where $\mathcal{L}_{\text{DAPT}}$ denotes the domain-adaptive pretraining loss and $\mathcal{L}_{QA}$ denotes the educational question–answer fine-tuning loss, with $\lambda$ controlling the emphasis on pedagogical alignment. In practice, these objectives are optimized sequentially rather than jointly: the model is first adapted to heliophysics literature using $\mathcal{L}_{\text{DAPT}}$, and the resulting DAPT-adapted checkpoint is subsequently fine-tuned using $\mathcal{L}_{QA}$ to produce the final SolarGPT-QA model. The combined objective summarizes the cumulative learning goal across training stages, rather than a single joint optimization. Unlike prior scientific language models optimized for expert-level analysis, SolarGPT-QA is explicitly designed for educational deployment. By combining domain-adaptive pretraining, parameter-efficient optimization, and pedagogically aligned fine-tuning with constrained decoding, the model bridges expert heliophysics knowledge and accessible K–12 instruction.
3.4 Inference time Hallucination Mitigation
In addition to domain-adaptive pretraining and educational fine-tuning, we applied several inference time controls to make responses more dependable. These methods do not modify the model’s parameters; instead, they guide decoding so that answers remain short, focused, and well formed. We used controlled sampling ($T=0.7$, top-$k=50$, top-$p=0.9$) with a repetition penalty of 1.1 to reduce repetitive continuations. We also restricted response length with a 128-token limit and a custom stopping criterion that ends generation at the first newline. Finally, we filtered the generated text to retain only complete sentences, removing unfinished fragments that could otherwise introduce vague content. These steps help limit hallucinations and improve answer clarity, as reflected in the ablation study in Section 4.6, where the models without inference time hallucination mitigation showed weaker results.
4 Experiments
4.1 Training Details and Compute